OpenStudy (anonymous):

Simplify (3a2 + 8a − 4) + (5a2 − 4).

OpenStudy (unklerhaukus):

\[(3a^2 + 8a − 4) + (5a^2 − 4)\]\[=3a^2 + 8a − 4+ 5a^2 − 4\]\[=3a^2 + 5a^2 + 8a − 4+− 4\]\[=(\cdots)a^2+(\cdots)a+(\cdots)\]